# Markovian dynamics for a quantum/classical system and quantum trajectories

### v1

@@ 2.4 Probabilities and instruments
A key point in the trajectory formulation is that the physical probabilities of Sec. 2.3.1 can be expressed by means of instruments, which are linear maps (see Def. 1); in this way the linearity of a quantum theory is respected. The SME (19) can not include arbitrary non-linearities: it is non-linear only due to normalization, and it is connected to a linear SME (12).
2.4.1 The stochastic map
By varying the initial condition in (12), we have that the linear SME defines a linear map on the trace class. As in Proposition 1, $X^{x}(t)$ is the solution of the stochastic equation (2) starting at $x$. Then, we consider the linear SME with the process $X^{x}(t)$ inside its coefficients (2.2); so, the random coefficients $\mathcal{L}_{t}$, $L_{kt}$, $J_{t}(u)$ in (12) are fixed and the equation is linear in the initial quantum state.
@@ 2.4.1 The stochastic map
$\Lambda^{t_{1},x}_{t_{2}}\circ\Lambda^{t_{2},x}_{t_{3}},\qquad 0\leq t_{1}\leq t_{2}\leq t_{3}.$ (23)
@@
The first properties and the composition rule (23) follow from the construction of $\sigma_{t}$ in Sec. 2.3 and from (12). By the fact that $X$ is a Markov process under $\mathbb{Q}$ and that this process appears in the right hand side of (12) only as $X(t_{-})$, the first property in (24) follows. The second property in (24) follows from (14a). ∎
@@ 2.4.2 The transition instruments
$\langle\rho,\mathcal{I}_{t}(\mathsf{E}|x)[a]\rangle=\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\langle\Lambda^{0,x}_{t\;*}[\rho],a\rangle\mathtt{1}_{\mathsf{E}}\big{(}X^{x}(t)\big{)}\right],\qquad\forall\rho\in\mathscr{T}(\mathscr{H}),\quad\forall a\in\mathscr{B}(\mathscr{H}),\quad\forall\mathsf{E}\in\mathcal{B}(\mathbb{R}^{s}).$ (26)
@@
By construction, equation (26) defines an instrument $\mathcal{I}_{t}(\cdot|x)$ on the $\sigma$-algebra $\mathcal{B}(\mathbb{R}^{s})$ (see Def. 1). Note that
@@
Moreover, the family of instruments (26) enjoys the following composition property, a quantum analogue of the Chapman-Kolmogorov identity for transition probabilities (57):
@@
Indeed, by using (23) and (24), we have
@@
$\langle\rho,\mathcal{I}_{t+t^{\prime}}(\mathsf{E}|x)[a]\rangle=\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\langle\Lambda^{0,x}_{t^{\prime}\;*}[\rho],\;\Lambda^{t^{\prime},x}_{t^{\prime}+t}[a]\rangle\mathtt{1}_{\mathsf{E}}\big{(}X^{x}(t+t^{\prime})\big{)}\Big{|}\mathscr{E}_{t^{\prime}}\right]\right]\\ {}=\int_{z\in\mathbb{R}^{s}}\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\langle\Lambda^{0,x}_{t^{\prime}\;*}[\rho],\;\Lambda^{t^{\prime},x}_{t^{\prime}+t}[a]\rangle\mathtt{1}_{\mathsf{E}-z}\big{(}X^{x}(t+t^{\prime})-X^{x}(t^{\prime})\big{)}\mathtt{1}_{\mathrm{d}z}\big{(}X^{x}(t^{\prime})\big{)}\Big{|}\mathscr{E}_{t^{\prime}}\right]\right]\\ {}=\int_{z\in\mathbb{R}^{s}}\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\langle\Lambda^{0,x}_{t^{\prime}\;*}[\rho],\;\mathcal{I}_{t}(E|z)[a]\rangle\mathtt{1}_{\mathrm{d}z}\big{(}X^{x}(t^{\prime})\big{)}\right]=\int_{z\in\mathbb{R}^{s}}\langle\rho,\;\mathcal{I}_{t^{\prime}}(\mathrm{d}z|x)\circ\mathcal{I}_{t}(\mathsf{E}|z)[a]\rangle.$
@@
The transition instruments (26) are uniquely determined by their Fourier transform, the characteristic operator
@@
$\langle\rho,\mathcal{G}_{t}(k|x)[a]\rangle=\int_{\mathbb{R}^{s}}\mathrm{e}^{\mathrm{i}k\cdot y}\langle\rho,\mathcal{I}_{t}(\mathrm{d}y|x)[a]\rangle=\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\langle\Lambda^{0,x}_{t\;*}[\rho],a\rangle\exp\left\{\mathrm{i}k\cdot X^{x}(t)\right\}\right],\qquad k\in\mathbb{R}^{s}.$ (29)
@@
The “Markovian” character of the hybrid dynamics is expressed by the fact that the instruments contain only the length of the time interval (not initial and final times) and that equation (28) holds: the transition instruments are “time homogeneous”, as the classical transition functions for a time homogeneous Markov process, compare the properties of instruments and equations (26)–(28) with equations (56)–(57) in Appendix A.
@@
Property (28) represents a compatibility condition among the various instruments at different times. Via Kolmogorov’s extension theorem [40, Theor. 1.8], this property allows to reconstruct the whole probability law of the process $X(t)$ starting from the joint probabilities at the times $0\leq t_{1}<t_{2}<\cdots<t_{m}$, given by
@@
$\mathbb{P}[X(t_{1})\in\mathsf{E}_{1},X(t_{2})\in\mathsf{E}_{2},\ldots,X(t_{m})\in\mathsf{E}_{m}]=\operatorname{\mathbb{E}}_{\mathbb{Q}}\big{[}\langle\sigma_{t_{m}},\mathds{1}\rangle\mathtt{1}_{\mathsf{E}_{1}}\big{(}X(t_{1})\big{)}\mathtt{1}_{\mathsf{E}_{2}}\big{(}X(t_{2})\big{)}\cdots\mathtt{1}_{\mathsf{E}_{m}}\big{(}X(t_{m})\big{)}\big{]}\\ {}=\int_{\mathbb{R}^{s}}\mathrm{d}x\big{\langle}\sigma_{0}(x),\;\int_{\mathsf{E}_{1}}\mathcal{I}_{t_{1}}(\mathrm{d}x_{1}|x)\circ\int_{\mathsf{E}_{2}}\mathcal{I}_{t_{2}-t_{1}}(\mathrm{d}x_{2}|x_{1})\circ\cdots\circ\int_{\mathsf{E}_{m}}\mathcal{I}_{t_{m}-t_{m-1}}(\mathrm{d}x_{m}|x_{m-1})[\mathds{1}]\big{\rangle},$ (30)
$\sigma_{0}(x)=\operatorname{\mathbb{E}}_{\mathbb{Q}}\big{[}\sigma_{0}\big{|}X(0)=x\big{]};$
here $\mathbb{P}$ is the probability defined in Sec. 2.3.1, which depends on the initial conditions $\sigma_{0}$, $X(0)$.
@@ 3 Quantum-classical dynamical semigroup
The hybrid dynamics has been introduced through the formalism of quantum trajectories; however, a Markovian hybrid dynamics is usually introduced by proposing suitable master equations. The most standard choice is to take probability densities (with respect to Lebesgue measure) as state space of the classical component [1, 2, 4, 3, 6, 14], but this choice is too restrictive in our case, as shown in Appendix A for the pure classical case.
A more general approach is to construct a dynamical semigroup in a suitable “space of observables” [7, 13] and to take for the classical component the spaces of functions $C_{\rm b}(\mathbb{R}^{s})$ and $B_{\rm b}(\mathbb{R}^{s})$, introduced in Sec. 1.1. Then, for the hybrid system we introduce the $C^{*}$-algebras [7]: margin: $\mathscr{A}_{1}$ margin: $\mathscr{A}_{2}$
@@
The elements of $\mathscr{A}_{2}$ are bounded functions from $\mathbb{R}^{s}$ to $\mathscr{B}(\mathscr{H})$, while $\mathscr{A}_{1}\subset\mathscr{A}_{2}$ is made up by the continuous bounded functions. The norm of $F\in\mathscr{A}_{i}$ is given by $\left\|F\right\|=\sup_{x\in\mathbb{R}^{s}}\left\|F(x)\right\|$, where $\left\|F(x)\right\|$ is the norm in $\mathscr{B}(\mathscr{H})$. The unit element $\mathbb{I}$ in $\mathcal{A}_{i}$ is given by $\mathbb{I}=\mathds{1}\otimes 1$;margin: $\mathbb{I}$ we can also write $\mathbb{I}(x)=\mathds{1}$.
@@ 3.1 The hybrid semigroup
$\langle\rho,\mathcal{I}_{t}(\mathsf{E}|x)[a]\rangle=\langle\rho,\mathcal{T}_{t}[a\otimes\mathtt{1}_{\mathsf{E}}](x)\rangle,\qquad\langle\rho,\mathcal{G}_{t}(k|x)[a]\rangle=\langle\rho,\mathcal{T}_{t}[a\otimes f_{k}]\rangle,\quad f_{k}(x)=\mathrm{e}^{\mathrm{i}k\cdot x}.$ (33)
To get the characteristic operators it is enough to have $\mathcal{T}_{t}$ defined on $\mathscr{A}_{1}$, while to get directly the instruments we need $\mathcal{T}_{t}$ defined on $\mathscr{A}_{2}$. As $\mathcal{T}_{t}$, $t\geq 0$, turns out to be a semigroup we call it the hybrid dynamical semigroup associated with the stochastic dynamics (2), (12).
@@
By 2-positivity, linearity, and the preservation of the unit, one gets a Kadison-like inequality [22, Sec. 3.1.1]; [44, Ex. 6.7 p. 107]:
@@
Under these conditions the associated dynamical semigroup (32) has been fully studied under the name of convolution semigroup of instruments and generalized to cases with a classical component not only in $\mathbb{R}^{s}$ [45, 41, 46, 47, 48, 49, 50]. In this construction a key point has been to exploit the analogies with the classical infinite divisible distributions [40, 39].
@@ 3.2 The generator
$\mathcal{K}[a\otimes f](x)=f(x)\mathcal{L}(x)[a]+a\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,c_{i}(x)+a\sum_{i,j=1}^{s}\frac{1}{2}\,\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}a_{i}^{k}(x)a_{j}^{k}(x)\\ {}+\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\sum_{k=1}^{d}a_{i}^{k}(x)\left(aL_{k}(x)+L_{k}(x)^{\dagger}a\right)\\ {}+\int_{\mathscr{U}}\left(\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\bigl{(}J(x,u)^{\dagger}+\mathds{1}\bigr{)}a\bigl{(}J(x,u)+\mathds{1}\bigr{)}-a\mathtt{1}_{\mathsf{U}_{0}}(u)\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(x,u)\right)\nu(\mathrm{d}u);$ (38)
$\mathcal{L}(x)$ is defined in (13b).
@@
$\mathrm{d}\left(f\big{(}X(t)\big{)}\langle\sigma_{t},a\rangle\right)\Big{|}_{X(t_{-})=x}=f(x)\biggl{(}\langle\sigma_{t_{-}},\mathcal{L}(x)[a]\rangle\mathrm{d}t+\sum_{k=1}^{d}\langle\sigma_{t_{-}},{L_{k}(x)}^{\dagger}a+aL_{k}(x)\rangle\mathrm{d}W_{k}(t)\\ {}+\int_{\mathscr{U}}\langle\sigma_{t_{-}},J(x,u)^{\dagger}aJ(x,u)+J(x,u)^{\dagger}a+aJ(x,u)\rangle\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)\biggr{)}\\ {}+\langle\sigma_{t_{-}},a\rangle\Biggl{(}\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\left(c_{i}(x)\mathrm{d}t+\sum_{k=1}^{d}a_{i}^{k}(x)\mathrm{d}W_{k}(t)\right)+\frac{1}{2}\sum_{i,j=1}^{s}\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}a_{i}^{k}(x)a_{j}^{k}(x)\mathrm{d}t\\ {}+\mathrm{d}t\int_{\mathsf{U}_{0}}\left(f\big{(}x+g(x,u)\big{)}-f(x)-\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(x,u)\right)\nu(\mathrm{d}u)\\ {}+\int_{\mathsf{U}_{0}}\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)+\int_{\mathscr{U}\setminus\mathsf{U}_{0}}\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\Pi(\mathrm{d}u,\mathrm{d}t)\Biggr{)}\\ {}+\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\sum_{k=1}^{d}a_{i}^{k}(x)\langle\sigma_{t_{-}},\,aL_{k}(x)+L_{k}(x)^{\dagger}a\rangle\mathrm{d}t\\ {}+\int_{\mathscr{U}}\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\langle\sigma_{t_{-}},J(x,u)^{\dagger}aJ(x,u)+J(x,u)^{\dagger}a+aJ(x,u)\rangle\Pi(\mathrm{d}u,\mathrm{d}t).$
@@
$\mathcal{K}[a\otimes f](x)=f(x)\mathcal{L}(x)[a]+a\Biggl{(}\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,c_{i}(x)+\frac{1}{2}\sum_{i,j=1}^{s}\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}a_{i}^{k}(x)a_{j}^{k}(x)\\ {}+\int_{\mathscr{U}}\left(f\big{(}x+g(x,u)\big{)}-f(x)-\mathtt{1}_{\mathsf{U}_{0}}(u)\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(x,u)\right)\nu(\mathrm{d}u)\Biggr{)}\\ {}+\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\sum_{k=1}^{d}a_{i}^{k}(x)\left(aL_{k}(x)+L_{k}(x)^{\dagger}a\right)\\ {}+\int_{\mathscr{U}}\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\left(J(x,u)^{\dagger}aJ(x,u)+J(x,u)^{\dagger}a+aJ(x,u)\right)\nu(\mathrm{d}u),$
@@
Here we have not studied the problem of determining the dynamical semigroup $\mathcal{T}_{t}$ from the generator (38). The dynamics of the hybrid system is fully determined by the SDEs (2) and (12) under Assumptions 1, 2, 3 (see Proposition 1, Theorem 3). To have the explicit expression (38) of the formal generator could help in clarifying the various physical interactions involved in the hybrid dynamics and it allows to compare the trajectory approach with the approaches based on master equations.

### v2

@@ 2.4 Probabilities and instruments
A key point in the trajectory formulation is that the physical probabilities of Sec. 2.3.1 can be expressed by means of instruments, which are completely positive linear maps (see Def. 1); in this way the linearity of a quantum theory is respected. The SME (22) can not include arbitrary non-linearities: it is non-linear only due to normalization and the change of probability; moreover, it is connected to the linear SME (10).
2.4.1 The stochastic map.
By varying the initial condition in (10), we have that the linear SME defines a linear map on the trace class. As in Proposition 1, $X^{x}(t)$ is the solution of the stochastic equation (2) starting at $x$. Then, we consider the linear SME with the process $X^{x}(t)$ inside its coefficients (6); so, the random coefficients $\mathcal{L}_{t}$, $L_{kt}$, $J_{t}(u)$ in (10) are fixed and the equation is linear in the initial quantum state.
@@
$\Lambda^{t_{1},x}_{t_{2}}\circ\Lambda^{t_{2},x}_{t_{3}}=\Lambda^{t_{1},x}_{t_{3}},\qquad 0\leq t_{1}\leq t_{2}\leq t_{3}.$ (26)
@@
The first properties and the composition rule (26) follow from the construction of $\sigma_{t}$ in Sec. 2.3 and from (10). By the fact that $X$ is a Markov process under $\mathbb{Q}$ and that this process appears in the right hand side of (10) only as $X(t_{-})$, the first property in (27) follows. The second property in (27) follows from (13). ∎
@@
$\langle\rho,\mathcal{I}_{t}(\mathsf{E}|x)[a]\rangle=\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\langle\Lambda^{0,x}_{t\;*}[\rho],a\rangle\mathtt{1}_{\mathsf{E}}\big{(}X^{x}(t)\big{)}\right],\qquad\forall\rho\in\mathscr{T}(\mathscr{H}),\qquad\forall a\in\mathscr{B}(\mathscr{H}),\qquad\forall\mathsf{E}\in\mathcal{B}(\mathbb{R}^{s}).$ (29)
@@
Moreover, we shall denote by $\delta_{x}(\cdot)$ margin: $\delta_{x}(\cdot)$ the trivial measure concentrated in the point $x\in\mathbb{R}^{s}$ (Dirac measure).
Proposition 5.
Equation (29) defines an instrument $\mathcal{I}_{t}(\cdot|x)$ on the $\sigma$-algebra $\mathcal{B}(\mathbb{R}^{s})$ (see Def. 1). For $t=0$ we get the trivial instrument
@@
Moreover, the family of instruments (29) enjoys the following composition property
@@
Equation (31) is a quantum analogue of the Chapman-Kolmogorov identity for transition probabilities (64).
Proof.
By using the properties of the maps $\Lambda^{0,x}_{t}$, given in Proposition 4, it is possible to check that the maps defined in (29) satisfy all the properties needed in the definition of instrument (Def. 1).
The composition property (31) is proved by direct computations. By using (26) and (27), we have
@@
$\langle\rho,\mathcal{I}_{t+t^{\prime}}(\mathsf{E}|x)[a]\rangle=\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\langle\Lambda^{0,x}_{t^{\prime}\;*}[\rho],\;\Lambda^{t^{\prime},x}_{t^{\prime}+t}[a]\rangle\mathtt{1}_{\mathsf{E}}\big{(}X^{x}(t+t^{\prime})\big{)}\Big{|}\mathscr{E}_{t^{\prime}}\right]\right]\\ {}=\int_{z\in\mathbb{R}^{s}}\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\langle\Lambda^{0,x}_{t^{\prime}\;*}[\rho],\;\Lambda^{t^{\prime},x}_{t^{\prime}+t}[a]\rangle\mathtt{1}_{\mathsf{E}-z}\big{(}X^{x}(t+t^{\prime})-X^{x}(t^{\prime})\big{)}\mathtt{1}_{\mathrm{d}z}\big{(}X^{x}(t^{\prime})\big{)}\Big{|}\mathscr{E}_{t^{\prime}}\right]\right]\\ {}=\int_{z\in\mathbb{R}^{s}}\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\langle\Lambda^{0,x}_{t^{\prime}\;*}[\rho],\;\mathcal{I}_{t}(E|z)[a]\rangle\mathtt{1}_{\mathrm{d}z}\big{(}X^{x}(t^{\prime})\big{)}\right]=\int_{z\in\mathbb{R}^{s}}\langle\rho,\;\mathcal{I}_{t^{\prime}}(\mathrm{d}z|x)\circ\mathcal{I}_{t}(\mathsf{E}|z)[a]\rangle;$
$\mathtt{1}_{\mathrm{d}z}\big{(}X^{x}(t^{\prime})\big{)}$ is the measure-theoretic notation for $\delta\big{(}z-X^{x}(t^{\prime})\big{)}\mathrm{d}z$; the quantity $X^{x}(t^{\prime})$ is implicitly contained in $\Lambda^{0,x}_{t^{\prime}\;*}$. ∎
@@
The transition instruments (29) are uniquely determined by their Fourier transform, the characteristic operator: for $k\in\mathbb{R}^{s}$,
@@
$\langle\rho,\mathcal{G}_{t}(k|x)[a]\rangle=\int_{\mathbb{R}^{s}}\mathrm{e}^{\mathrm{i}k\cdot y}\langle\rho,\mathcal{I}_{t}(\mathrm{d}y|x)[a]\rangle=\operatorname{\mathbb{E}}_{\mathbb{Q}}\left[\langle\Lambda^{0,x}_{t\;*}[\rho],a\rangle\exp\left\{\mathrm{i}k\cdot X^{x}(t)\right\}\right].$ (32)
@@
The “Markovian” character of the hybrid dynamics is expressed by the fact that the instruments contain only the length of the time interval (not initial and final times) and that equation (31) holds: the transition instruments are “time-homogeneous”, as the classical transition functions for a time-homogeneous Markov process, compare the properties of instruments and equations (29)–(31) with equations (63)–(64).
@@
Property (31) represents a compatibility condition among the various instruments at different times. Via Kolmogorov’s extension theorem [45, Theor. 1.8], this property allows to reconstruct the whole probability law of the process $X(t)$ starting from the joint probabilities at the times $0\leq t_{1}<t_{2}<\cdots<t_{m}$, given by
@@
$\mathbb{P}[X(t_{1})\in\mathsf{E}_{1},X(t_{2})\in\mathsf{E}_{2},\ldots,X(t_{m})\in\mathsf{E}_{m}]=\operatorname{\mathbb{E}}_{\mathbb{Q}}\big{[}\langle\sigma_{t_{m}},\mathds{1}\rangle\mathtt{1}_{\mathsf{E}_{1}}\big{(}X(t_{1})\big{)}\mathtt{1}_{\mathsf{E}_{2}}\big{(}X(t_{2})\big{)}\cdots\mathtt{1}_{\mathsf{E}_{m}}\big{(}X(t_{m})\big{)}\big{]}\\ {}=\int_{\mathbb{R}^{s}}\mathbb{Q}_{0}(\mathrm{d}x)\big{\langle}\sigma_{0}(x),\;\int_{\mathsf{E}_{1}}\mathcal{I}_{t_{1}}(\mathrm{d}x_{1}|x)\circ\int_{\mathsf{E}_{2}}\mathcal{I}_{t_{2}-t_{1}}(\mathrm{d}x_{2}|x_{1})\circ\cdots\circ\int_{\mathsf{E}_{m}}\mathcal{I}_{t_{m}-t_{m-1}}(\mathrm{d}x_{m}|x_{m-1})[\mathds{1}]\big{\rangle},$ (33)
$\sigma_{0}(x)=\operatorname{\mathbb{E}}_{\mathbb{Q}}\big{[}\sigma_{0}\big{|}X(0)=x\big{]},\qquad\mathbb{Q}_{0}(\mathsf{E}):=\mathbb{Q}[X(0)\in\mathsf{E}];$
here $\mathbb{P}$ is the probability defined in Sec. 2.3.1, which depends on the initial conditions $\sigma_{0}$, $X(0)$. Note that we have $\langle\sigma_{t_{m}},\mathds{1}\rangle=\operatorname{Tr}\{\sigma_{t_{m}}\}=p_{t_{m}}$, which is the probability density for events up to the biggest time $t_{m}$, in agreement with the discussion following the definition (17) of $\mathbb{P}$.
In the construction of the hybrid dynamics, we have used a two-step procedure: firstly, the SDEs (2) and (10) under the reference probability $\mathbb{Q}$ have been introduced, and, then, the SDEs (20) and (22) under the physical probability $\mathbb{P}$. By following this path, already developed in the field of time-continuous measurements [23], we were able to give rise to the instruments (29) with the right properties, given in Proposition 5, and with the connection (33) with the physical probability $\mathbb{P}$. In this way, the connection with the general formulation of a quantum theory is preserved.
@@ 3 Quantum-classical dynamical semigroup
The hybrid dynamics has been introduced through the formalism of quantum trajectories; however, a Markovian hybrid dynamics is usually introduced by proposing suitable master equations. The most standard choice is to take probability densities (with respect to Lebesgue measure) as state space of the classical component [1, 2, 4, 3, 6, 14, 5], but this choice is too restrictive in our case, as shown in A for the pure classical case.
A more general approach is to construct a dynamical semigroup in a suitable “space of observables” [7, 13] and to take for the classical component the spaces of functions $C_{\rm b}(\mathbb{R}^{s})$ and $B_{\rm b}(\mathbb{R}^{s})$, introduced in Sec. 1.1. Then, for the hybrid system we introduce margin: $\mathscr{A}_{1}$ the $C^{*}$-algebras [7]:
@@
The elements of $\mathscr{A}_{2}$ margin: $\mathscr{A}_{2}$ are bounded functions from $\mathbb{R}^{s}$ to $\mathscr{B}(\mathscr{H})$, while $\mathscr{A}_{1}\subset\mathscr{A}_{2}$ is made up by the continuous bounded functions. The norm of $F\in\mathscr{A}_{i}$ is given by $\left\|F\right\|=\sup_{x\in\mathbb{R}^{s}}\left\|F(x)\right\|$, where $\left\|F(x)\right\|$ is the norm in $\mathscr{B}(\mathscr{H})$. The unit element $\mathbb{I}$ in $\mathcal{A}_{i}$ is given by $\mathbb{I}=\mathds{1}\otimes 1$;margin: $\mathbb{I}$ we can also write $\mathbb{I}(x)=\mathds{1}$.
@@ 3.1 The hybrid semigroup
$\langle\rho,\mathcal{I}_{t}(\mathsf{E}|x)[a]\rangle=\langle\rho,\mathcal{T}_{t}[a\otimes\mathtt{1}_{\mathsf{E}}](x)\rangle,\quad\langle\rho,\mathcal{G}_{t}(k|x)[a]\rangle=\langle\rho,\mathcal{T}_{t}[a\otimes f_{k}]\rangle,\quad f_{k}(x)=\mathrm{e}^{\mathrm{i}k\cdot x}.$ (36)
To get the characteristic operators it is enough to have $\mathcal{T}_{t}$ defined on $\mathscr{A}_{1}$, while to get directly the instruments we need $\mathcal{T}_{t}$ defined on $\mathscr{A}_{2}$. As $\mathcal{T}_{t}$, $t\geq 0$, turns out to be a semigroup, we call it the hybrid dynamical semigroup associated with the stochastic dynamics (2), (10).
@@
By 2-positivity, linearity, and the preservation of the unit, one gets a Kadison-like inequality [22, Sec. 3.1.1], [50, Ex. 6.7 p. 107]:
@@
Under these conditions the associated dynamical semigroup (35) has been fully studied under the name of convolution semigroup of instruments and generalized to cases with a classical component not only in $\mathbb{R}^{s}$ [51, 46, 52, 53, 54, 55, 56, 27]. In this construction a key point has been to exploit the analogies with the classical infinitely divisible distributions [45, 44].
The dual $C^{*}$-algebra of $\mathscr{A}_{2}$ is the space containing the hybrid states and the adjoint semigroup ${\mathcal{T}_{t}}^{*}$ gives the dynamics of the hybrid states; therefore, the hybrid dynamics is completely positive and linear. Another relevant point is Eq. (36), which defines the transition instruments $\mathcal{I}_{t}(\cdot|x)$ once a semigroup $\mathcal{T}_{t}$, satisfying the properties in Proposition 6, is given. Therefore, the dynamics ${\mathcal{T}_{t}}^{*}$ gives not only the hybrid state at the time $t$, but also all the multi-time probabilities (33). Once again, to give a hybrid dynamical semigroup (linear and completely positive) allows to respect the general structure of a quantum theory, as discussed at the end of Sec. 2.4.2 in the case of the stochastic formulation.
@@ 3.2 The generator
$\mathcal{K}[a\otimes f](x)=f(x)\mathcal{L}(x)[a]+a\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,c_{i}(x)+a\sum_{i,j=1}^{s}\frac{1}{2}\,\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}b_{i}^{k}(x)b_{j}^{k}(x)\\ +\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\sum_{k=1}^{d}b_{i}^{k}(x)\left(aL_{k}(x)+L_{k}(x)^{\dagger}a\right)\\ +\int_{\mathscr{U}}\biggr{(}\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\bigl{(}J(x,u)^{\dagger}+\mathds{1}\bigr{)}a\bigl{(}J(x,u)+\mathds{1}\bigr{)}-a\mathtt{1}_{\mathsf{U}_{0}}(u)\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(x,u)\biggl{)}\nu(\mathrm{d}u);$ (41)
$\mathcal{L}(x)$ is defined in (12).
@@
$\mathrm{d}\left(f\big{(}X(t)\big{)}\langle\sigma_{t},a\rangle\right)\Big{|}_{X(t_{-})=x}=f(x)\biggl{(}\langle\sigma_{t_{-}},\mathcal{L}(x)[a]\rangle\mathrm{d}t+\sum_{k=1}^{d}\langle\sigma_{t_{-}},{L_{k}(x)}^{\dagger}a+aL_{k}(x)\rangle\mathrm{d}W_{k}(t)\\ {}+\int_{\mathscr{U}}\langle\sigma_{t_{-}},J(x,u)^{\dagger}aJ(x,u)+J(x,u)^{\dagger}a+aJ(x,u)\rangle\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)\biggr{)}\\ {}+\langle\sigma_{t_{-}},a\rangle\Biggl{(}\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\left(c_{i}(x)\mathrm{d}t+\sum_{k=1}^{d}b_{i}^{k}(x)\mathrm{d}W_{k}(t)\right)+\frac{1}{2}\sum_{i,j=1}^{s}\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}b_{i}^{k}(x)b_{j}^{k}(x)\mathrm{d}t\\ {}+\mathrm{d}t\int_{\mathsf{U}_{0}}\left(f\big{(}x+g(x,u)\big{)}-f(x)-\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(x,u)\right)\nu(\mathrm{d}u)\\ {}+\int_{\mathsf{U}_{0}}\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)+\int_{\mathscr{U}\setminus\mathsf{U}_{0}}\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\Pi(\mathrm{d}u,\mathrm{d}t)\Biggr{)}\\ {}+\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\sum_{k=1}^{d}b_{i}^{k}(x)\langle\sigma_{t_{-}},\,aL_{k}(x)+L_{k}(x)^{\dagger}a\rangle\mathrm{d}t\\ {}+\int_{\mathscr{U}}\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\langle\sigma_{t_{-}},J(x,u)^{\dagger}aJ(x,u)+J(x,u)^{\dagger}a+aJ(x,u)\rangle\Pi(\mathrm{d}u,\mathrm{d}t).$
@@
$\mathcal{K}[a\otimes f](x)=f(x)\mathcal{L}(x)[a]+a\Biggl{(}\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,c_{i}(x)+\frac{1}{2}\sum_{i,j=1}^{s}\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}b_{i}^{k}(x)b_{j}^{k}(x)\\ {}+\int_{\mathscr{U}}\left(f\big{(}x+g(x,u)\big{)}-f(x)-\mathtt{1}_{\mathsf{U}_{0}}(u)\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(x,u)\right)\nu(\mathrm{d}u)\Biggr{)}\\ {}+\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\sum_{k=1}^{d}b_{i}^{k}(x)\left(aL_{k}(x)+L_{k}(x)^{\dagger}a\right)\\ {}\quad{}+\int_{\mathscr{U}}\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\left(J(x,u)^{\dagger}aJ(x,u)+J(x,u)^{\dagger}a+aJ(x,u)\right)\nu(\mathrm{d}u),$
@@
Here we have not studied the problem of determining the dynamical semigroup $\mathcal{T}_{t}$ from the generator (41). By construction $\mathcal{T}_{t}$, defined by (34), is a solution of (40); what is not proved is the uniqueness of this solution. The dynamics of the hybrid system is fully determined by the stochastic differential equations (2) and (10) under Assumptions 1, 2, 3 (see Proposition 1, Theorem 3).